### v1

# NaviAgent: Bilevel Planning on Tool Dependency Graphs for Function Calling

Abstract
While large language models (LLMs) have demonstrated remarkable potential in open-domain reasoning, their reliance on static knowledge and fragile tool invocation severely hinder the orchestration of complex, heterogeneous toolchains, especially when scaled to thousands of tools. Existing methods typically use rigid single-path execution, resulting in poor error recovery and exponentially growing search spaces. We introduce NaviAgent, a graph-navigated bilevel planning architecture for robust function calling, comprising a Multi-Path Decider and Graph-Encoded Navigator. As an LLM-powered agent, the Multi-Path Decider defines a four-dimensional decision space and continuously perceives environmental states, dynamically selecting the optimal action to fully cover all tool invocation scenarios. The Graph-Encoded Navigator constructs a Tool Dependency Heterogeneous Graph (TDHG), where node embeddings explicitly fuse API schema structure with historical invocation behavior. It also integrates a novel heuristic search strategy that guides the Decider toward efficient and highly successful toolchains, even for unseen tool combinations. Experiments show that NaviAgent consistently achieves the highest task success rate (TSR) across all foundation models and task complexities, outperforming the average baselines (ReAct, ToolLLM, $\alpha$-UMI) by 13.5%, 16.4%, and 19.0% on Qwen2.5-14B, Qwen2.5-32B, and Deepseek-V3, respectively. Its execution steps are typically within one step of the most efficient baseline, ensuring a strong balance between quality and efficiency. Notably, a fine-tuned Qwen2.5-14B model achieves a TSR of 49.5%, surpassing the much larger 32B model (44.9%) under our architecture. Incorporating the Graph-Encoded Navigator further boosts TSR by an average of 2.4 points, with gains up over 9 points on complex tasks for larger models (Deepseek-V3 and GPT-4o), highlighting its essential role in toolchain orchestration.
1 Introduction
Integration of large language models (LLMs) with external tools has evolved from single-tool interactions parisi2022talm ; schick2023toolformer ; chen2024towards to function calling agents shen2023hugginggpt ; yang2023gpt4tools capable of executing complex, multi-stage procedures qu2025tool . However, critical real-world challenges such as API instability, rigid workflow specifications remain unresolved. Despite incremental improvements from Chameleon’s modular heterogeneous tool integration lu2023chameleon , ToolVerifier’s self-verification paradigm mekala2024toolverifier , and Confucius’s iterative self-guidance mechanism gao2024confucius , all such methods ultimately inherit the weakness of static workflow architectures maldonado2024multi ; krishnan2025ai - an inability to dynamically adapt to runtime contingencies. When critical API services are unavailable, they lack mechanisms for dynamic fallback tran2025multi ; jiang2025appagentx or alternative path discovery, severely limiting real-world deployment reliability, especially in complex multi-tool chaining call scenarios barua2024exploring ; shi2024chain .
Compounding these challenges, scalability issues escalate dramatically as tool libraries grow. Heuristic approaches like ToolLLM qin2023toolllm and ToolChain* zhuang2023toolchain reduce search complexity through functional hierarchy trees and A* cost estimation, yet remain fundamentally constrained by static tool relationship assumptions. While ToolNet liu2024toolnet attempts to address this via dynamic dependency graphs constructed from historical invocation logs, its practical utility is limited by sparse multi-hop interactions, only 12% of APIs in production logs exhibit more than 3 chained calls, leaving long-tail tools disconnected. This sparsity creates a cold-start bottleneck zhang2025cold ; zhang2024llmtreerec , raising a core question: Can we automatically infer potential tool collaboration patterns by combining inherent structural features (e.g., parameter types, I/O constraints) rather than relying solely on historical data?
Building upon the graph-enhanced LLM paradigm wu2024can , we present NaviAgent, a novel graph-guided framework for robust function calling. NaviAgent features a bilevel planning architecture that integrates a Multi-Path Decider and a Graph-Encoded Navigator, enabling dynamic toolchain orchestration and error recovery. The Multi-Path Decider defines a four-dimensional decision space and continuously perceives environmental states to flexibly select optimal actions for both routine and exceptional scenarios. The Graph-Encoded Navigator, with a novel heuristic search strategy, constructs and navigates a Tool Dependency Heterogeneous Graph (TDHG) that captures both structural and behavioral relationships among tools. This closed-loop mechanism integrates graph-based reasoning with real-time execution feedback, allowing the agent to discover alternative paths and adapt to runtime contingencies. As a result, NaviAgent robustly handles both simple and complex tool invocation scenarios, supporting scalable and reliable deployment in large, heterogeneous tool environments. Our experiments show that NaviAgent substantially outperforms the strongest baseline ($\alpha$-UMI) on Deepseek-V3, achieving improvements of 15.4% in task completion rate and 14.8% in task success rate (reaching 98.3% and 55.5%, respectively). The integration of the Graph-Encoded Navigator consistently provides a average 2.4 point boost over non-graph baselines across all tasks, with gains of up to 8 points on complex tasks for larger models, underscoring its critical role in complex toolchain coordination. Furthermore, our framework enables a fine-tuned lightweight Qwen2.5-14B model to even surpass the much larger Qwen2.5-32B model in task success rate (49.5% vs. 44.9%), demonstrating NaviAgent’s efficiency and scalability.
Our main contributions are as follows: i) NaviAgent Architecture: A scalable bilevel planning framework that orchestrates toolchains at thousand-scale while achieving superior performance to SOTA baselines, maintaining an optimal quality-efficiency equilibrium. ii) Multi-Path Decider: A four-dimensional decision space covering the full spectrum of tool invocation scenarios. iii) Dynamic Graph-Encoded Navigator: Joint learning of node features and adaptive edge weights enables continuous optimization and discovery of both existing and novel tool combinations.
2 Related Work
Single-Tool Invocation.
Early research focused on enhancing LLMs’ single-tool invocation capabilities. TALM parisi2022talm established foundational paradigms through predefined template chains, while EasyTool yuan2024easytool introduced structured tool descriptions to reduce semantic parsing overhead. For long-context scenarios, tool documentation compression techniques preserved critical semantics via summarization, enabling low-resource tool usage xu2024concise . Toolformer schick2023toolformer innovatively embedded tool invocation APIs in pre-training, allowing self-supervised learning of usage patterns from unlabeled data. In multimodal settings, GPT4Tools yang2023gpt4tools improved visual tool generalization (e.g. object detection) by aligning vision-language instructions with tool descriptions.
Multi-Tool Orchestration.
As tool libraries expanded, HuggingGPT shen2023hugginggpt proposed a four-stage pipeline (plan, select, execute, respond) for standardized multi-tool workflows, while Chameleon lu2023chameleon integrated heterogeneous tools (13+ types) via modular composition. Similarly, $\alpha$-UMI shen2024small decomposes the tool-use process into planning, invocation, and summarization, but uniquely assigns each stage to a dedicated lightweight LLM, enabling modular updates and improved performance, especially for smaller models. For small toolkits, TRICE qiao2023making optimized single tool policies via execution feedback, and ToolFactory ni2025toolfactory automated tool adaptation through domain-guided code synthesis. However, these approaches struggled with dynamic collaboration. For large-scale toolkits, Confucius gao2024confucius addressed combinatorial explosion via hierarchical tool classification, and ToolVerifier mekala2024toolverifier improved selection robustness through self-verification mechanisms.
Dynamic Planning & Adaptation.
Static frameworks faltered under open-domain task complexity, prompting dynamic decision mechanisms. ReAct yao2023react pioneered the decoupling of reasoning from tool calls through chain-of-thought planning. Building on this, Reflexion shinn2023reflexion enhanced error recovery by introducing iterative self-reflection, significantly improving fault tolerance in complex tasks. For long-horizon tasks, path search techniques became pivotal: Tree-of-Thoughts (ToT) yao2023tree formalized tool invocation as searchable reasoning trees with dynamic branching, while ToolLLM qin2023toolllm optimized search efficiency through functional hierarchy-guided DFS. ToolChain zhuang2023toolchain further advanced this by employing heuristic cost estimation to prioritize high-success-rate branches. Yet, these methods assumed static tool relationships, failing to adapt to API drift or cross-domain tasks. ControlLLM liu2024controlllm built static dependency graphs for task decomposition, whereas ToolNet liu2024toolnet dynamically updated tool relations from historical calls, both limited by sparse multi-hop interaction data. This gap motivates our dynamic Graph-Encoded Navigator that jointly models structural dependencies and behavioral adaptations to capture evolving tool relations, aligning with findings that graph learning enhance LLM planning wu2024can ; besta2024graph .
3 Methodology
Figure 1: The overview of NaviAgent.
Figure 1 presents the NaviAgent architecture, which integrates a Multi-Path Decider and a dynamically updating Graph-Encoded Navigator in a bilevel planning framework. The Decider, shown in the blue box, selects among actions such as direct response, intent clarification, tool retrieval, or tool call. The Navigator, visualized as five dark brown components, is responsible for planning toolchains over an evolving TDHG with learned edge weights, as depicted in the graph structure and its temporal evolution. As the Decider interacts with the environment, the graph structure and edge weights are continuously updated based on execution feedback, enabling adaptive and robust planning. The bottom of the figure demonstrates NaviAgent’s ability to learn more optimal routing strategies over time, as well as to flexibly adapt its execution path in response to API unavailability. Further details on the graph evolution mechanism are provided in Section 3.2.4.
3.1 Multi-Path Decider
3.1.1 LLM-Driven Autonomous Agent Architecture
[…]
$\mathcal{H}_{t}=\langle(o_{t-3},a_{t-3}),\ldots,(o_{t-1},a_{t-1})\rangle$ (1)
where a sliding window maintains the most recent three [^1] Our experiments demonstrate that utilizing the most recent three observation-action pairs achieves the best balance between accuracy and efficiency. observation-action pairs, capturing the Decider’s recent decision trajectory. The pruned tool dependency subgraph $\mathcal{G}_{t-1}^{\prime}=(V,E,W)$ is computed based on the Decider’s state at the previous time step $t-1$, where $V$ is the node set, $E$ is the edge set, and $W$ denotes the edge weights indicating dependency strengths. The overall decision function is then formulated as
$a_{t}=F(\mathcal{H}_{t},\mathcal{O}_{t},\mathcal{G}_{t-1}^{\prime})$ (2)
where $\mathcal{O}_{t}$ is the current observation, and $a_{t}\in\mathcal{A}$ is the action selected at time $t$.
3.1.2 Model Training
For supervised fine-tuning, we adopt the standard language modeling objective, computing the loss exclusively over the response or action generation segments. During training, the LLM-based decider receives as input the most recent historical state-action pairs $\mathcal{H}_{t}$, the current observation $\mathcal{O}_{t}$, and the pruned tool dependency subgraph $\mathcal{G}_{\mathrm{sub}}$. The model is trained to maximize the likelihood of the ground-truth action $a_{t}^{*}$ at step $t$, which is derived from high-quality, curated datasets (see Appendix E.2 for details):
$\mathcal{L}_{\mathrm{SFT}}=-\frac{1}{N}\sum_{i=1}^{N}\log p_{\theta}(a_{t}^{*}\mid\mathcal{H}_{t},\mathcal{O}_{t},\mathcal{G}_{\mathrm{sub}})$ (3)
where $N$ is the number of training samples and $p_{\theta}$ denotes the Decider’s predicted probability over the action space.
3.2 Graph-Encoded Navigator
3.2.1 Graph Construction and Representation
While tool standardization frameworks (e.g., Anthropic’s MCP) help normalize basic API metadata, challenges remain due to inconsistent parameter naming and undocumented tool dependencies. In our framework, each tool consists of one or more APIs. We address these issues by applying semantic similarity clustering to unify functionally equivalent parameters (see details in Appendix A).
Definition.
We construct a TDHG, a directed weighted graph $\mathcal{G}=(V,E,W)$ with API and parameter nodes. Edges include structural chains, defined by API schemas (e.g., parameter-to-API and API-to-parameter connections), as well as behavioral chains, derived from historical usage data (e.g., API-to-API and parameter-to-parameter dependencies). Each edge is assigned a statistical weight $\tilde{w}_{ij}$ reflecting empirical invocation patterns.
$\tilde{w}_{ij}=\frac{N(v_{i}\rightarrow v_{j})}{N(v_{j})}$ (4)
where $N(v_{i}\rightarrow v_{j})$ counts the number of successful invocations from $v_{i}$ to $v_{j}$, and $N(v_{j})$ is the total number of invocations involving $v_{j}$.
We formulate tool dependency discovery as a link prediction problem hamilton2017inductive ; zhang2018link ; zhou2020graph ; wu2024can . To model this, we employ a Heterogeneous Graph Transformer (HGT) that integrates node-level feature fusion, type-specific encoding, and relation-aware message passing. Each node is initialized with both semantic (BGE-based) and structural features (including invocation statistics and degree information), and projected into a unified embedding space. We stack two multi-head HGT layers to aggregate information from the 2-hop neighborhood. Notably, the attention mechanism incorporates a statistical edge weight $\tilde{w}_{uv}$ to reflect empirical call patterns:
$\alpha_{uv}^{(k,r)}={softmax}_{u\in\mathcal{N}_{r}(v)}\left(\frac{(\mathbf{W}_{Q}^{(k,r)}\mathbf{h}_{u}^{\prime})^{\top}(\mathbf{W}_{K}^{(k,r)}\mathbf{h}_{v}^{\prime})}{\sqrt{d_{k}}}+\mathbf{b}_{r}^{(k)}+\tilde{w}_{uv}\right)$ (5)
where $\mathcal{N}_{r}(v)$ denotes the set of neighbors of node $v$ under relation $r$, and $\mathbf{h}_{u}^{\prime}$, $\mathbf{h}_{v}^{\prime}$ are the type-specific encoded representations of nodes $u$ and $v$ (see Appendix B.1 for details). $\mathbf{W}_{Q}^{(k,r)}$ and $\mathbf{W}_{K}^{(k,r)}$ are the query and key projection matrices for head $k$ and relation $r$, $\mathbf{b}_{r}^{(k)}$ is an edge-type-specific bias, and $d_{k}=d/8$ is the dimension per head. Then the concatenated head outputs are projected to obtain the final node embeddings, which are then used for link prediction.
3.2.2 Graph Training Objective
The graph model is trained with a hybrid loss that combines cross-entropy $\mathcal{L}_{CE}$ and adaptive margin $\mathcal{L}_{\mathrm{Margin}}$ objectives (see Appendix B.2 for details), both leveraging edge weights $\tilde{w}_{uv}$ to capture graded dependencies. Specifically, statistical edge weights serve as soft labels in the cross-entropy loss and dynamically adjust the margin in the margin loss, enabling the model to focus on more critical dependencies. The final objective is a weighted sum of these two terms:
$\mathcal{L}=\mu_{t}\cdot\mathcal{L}_{CE}+(1-\mu_{t})\cdot\mathcal{L}_{\mathrm{Margin}}$ (6)
where $\mu_{t}=\mu_{0}\cdot\gamma^{t}$ is a decaying weight over training epochs $t$, with initial value $\mu_{0}$ and decay rate $\gamma\in(0,1)$. This curriculum-inspired strategy first emphasizes classification accuracy and then prioritizes representation discrimination, leading to both robust prediction and well-structured embeddings.
3.2.3 Graph Search
At inference time, the predicted link probabilities $p_{uv}$ are assigned as edge weights $w_{uv}$ in the TDHG, providing the foundation for graph search and toolchain planning. We implement two classes of search strategies.
Alpha-Beta Pruning.
This algorithm knuth1975analysis is adapted for backward search over the TDHG, using dynamic thresholding and path scoring to prune suboptimal toolchains. The core scoring function balances direct edge weights, target connectivity, and path depth:
$S_{uv}=\frac{w_{uv}+\mathbb{I}(u\rightarrow v_{t}^{\mathrm{api}})w_{u\rightarrow v_{t}^{\mathrm{api}}}+\mathbb{I}(u\rightarrow v_{t}^{\mathrm{param}})w_{u\rightarrow v_{t}^{\mathrm{param}}}}{3}\times\mathcal{D}(d)$ (7)
where $w_{u\rightarrow v_{t}^{\mathrm{api}}}$ and $w_{u\rightarrow v_{t}^{\mathrm{param}}}$ denote the edge weights from $u$ to the target API node $v_{t}^{\mathrm{api}}$ and target parameter node $v_{t}^{\mathrm{param}}$, respectively, included only if the corresponding indicator function $\mathbb{I}(\cdot)$ is active. And $\mathcal{D}(d)$ penalizes longer paths.
Heuristic Graph Search with Dynamic Pruning.
Our hybrid heuristic search algorithm combines simulated annealing kirkpatrick1983optimization and genetic algorithm strategies shapiro1999genetic , using a composite fitness function to evaluate candidate toolchains:
$\mathcal{F}_{\omega}=0.35\mathcal{C}_{c}+0.15\log(1+\rho_{p})+0.3\mathcal{D}_{c}+0.15\mathcal{W}_{n}+0.05\mathcal{C}_{p}$ (8)
where $\mathcal{C}_{c}$ (node compactness) measures the closeness centrality of API nodes, $\rho_{p}$ (parameter density) is the ratio of parameter nodes within the subgraph to promote concise yet informative solutions, $\mathcal{D}_{c}=0.2e^{-d/10}+0.8e^{-n/8}$ (depth penalty) penalizes overly deep or complex dependency structures, with $d$ as the average depth and $n$ as the total node count, $\mathcal{W}_{n}$ (weight quantification) encourages solutions with higher cumulative edge weights, and $\mathcal{C}_{p}$ (path complexity) evaluates structural simplicity, favoring solutions with less intricate connectivity. Complete parameter settings, algorithmic details for both search strategies are provided in Appendix B.3 and Algorithm 1,  2).
3.2.4 Graph Evolution
The TDHG in real-world environments is inherently dynamic, evolving as new tools are introduced, obsolete ones are deprecated, and usage patterns shift. As shown at the bottom of Figure 1: At time $T-1$, selectable paths are relatively uniform (similar line shades), indicating multiple equally viable routes. By $T$, the Navigator has learned to prefer a more optimal path (the top, darker line). At $T+1$, when the upper-left API becomes unavailable, the Navigator adapts by recomputing and selecting the lower route as the new optimal path. This demonstrates the Navigator’s ability to learn from feedback and flexibly adjust its planning in response to runtime changes. To systematically support such adaptability, we design a graph evolution framework with three key mechanisms:
Incremental Node Integration.
To accommodate newly introduced tools, we incrementally add new nodes via semantic similarity clustering, initializing their parameters (e.g., $N_{succ}(v)=0$, $N_{fail}(v)=0$ for successful and failed invocation counts) and the statistical weights of associated edges (e.g., $\tilde{w}_{uv}=0$) to ensure consistency with existing graph features.
Targeted Subgraph Pruning.
Obsolete or rarely used tools are selectively pruned based on a weighted combination of failure rates and invocation frequencies:
$\mathrm{Prune}(v)\propto\lambda\cdot\sigma(f_{fail}(v))+(1-\lambda)\cdot\sigma(f_{freq}(v)^{-1})$ (9)
where $\lambda\in[0,1]$ controls the trade-off between failure rates and invocation frequencies, and $f_{fail}$ and $f_{freq}$ denote failure rates and invocation frequencies, respectively.
Edge Attribute Propagation.
Long-term stability and short-term adaptation are balanced by updating the statistical edge weights $\tilde{w}_{uv}$ through a combination of historical trends and recent invocation success rates:
$\tilde{w}_{uv}^{(t)}=\eta\cdot\underbrace{\tilde{w}_{uv}^{(t-1)}}_{\mathrm{long-term\ weight}}+(1-\eta)\cdot\underbrace{\frac{N_{succ}^{\mathrm{recent\ \tau\ days}}(u\to v)}{N_{succ}^{\mathrm{recent\ \tau\ days}}(v)}}_{\mathrm{recent\ success\ rate}}$ (10)
where $\eta\in[0,1]$ balances long-term memory and recent observations, and $N_{succ}^{\mathrm{recent\ \tau\ days}}$ denotes successful invocations within a sliding window of $\tau$ days. These dynamically updated statistical edge weights $\tilde{w}_{uv}$ are subsequently used as soft labels for supervising model training, as described in Section 3.2.2.
3.3 Dynamic Execution & Path Recombination
Robust and adaptive toolchain orchestration is achieved through a bilevel dynamic planning framework, in which the Decider manages action selection and the Navigator is responsible for toolchain planning.
Collaboration Workflow.
The case in Figure 2 demonstrates the workflow: given a multi-intent query, the Decider first decomposes it into two sub-tasks, one that can be answered directly and another that requires tool retrieval. For the sub-task requiring tool use, the Decider employs a BGE-based retriever to identify three target APIs (highlighted as red solid nodes), and invokes the Navigator to perform backward graph search, yielding three candidate subgraphs $\mathcal{G}^{\prime}$ (composed of red edges). These are then returned to the Decider, which executes the corresponding tool call. More detailed cases can be found in the Appendix C.
Figure 2: Case Study of the Collaboration Workflow.
Path Recombination.
If a tool call action fails, for example, due to an unavailable or malfunctioning API, the Decider will not terminate or rigidly adhere to the original plan. Instead, it selects the tool retrieval action, triggering a new round of collaboration with the Navigator. The Navigator then performs graph search on the current TDHG to identify an alternative toolchain. The Decider subsequently executes the new plan, and this process can be repeated as necessary until the task is completed or determined to be infeasible. This dynamic path recombination allows the agent to adaptively recover from failures and maximize task completion in complex tool environments.
4 Experiments
4.1 Experimental Settings
Datasets.
Our experiments are based on two public API benchmarks: API-Bank li2023api and ToolBench qin2023toolllm . As real-time API execution is currently unavailable, evaluation tasks are constructed in a simulated environment based on the extensive API lists and conversational trajectories provided by these datasets. Tasks are categorized into three levels of complexity: Easy (at most one API call or directly answerable), Medium (two API calls), and Hard (three or more APIs). Details of task generation are provided in Appendix D. For model fine-tuning, Qwen2.5-14B is trained on 3,500+ examples sampled from our generated task set, with strict separation between fine-tuning and evaluation data to prevent leakage.
Baselines and Models.
We compare our approach with several representative baselines, including ReAct yao2023react (alternates reasoning steps and tool calls), ToolLLM qin2023toolllm (DFSDT-based tool planning and execution), and $\alpha$-UMI shen2024small (multi-agent framework with modular tool-use stages handled by lightweight LLMs). Experiments are conducted across multiple foundation models, including open-source models (Qwen2.5-14B yang2024qwen2 , Qwen2.5-32B tahmid2024qwen2 , DeepSeek-R1-Distill-Qwen-32B(DeepSeek-R1-32B) guo2025deepseek ) and closed-source models (DeepSeek-V3 liu2024deepseek , GPT-4o hurst2024gpt ), as well as a fine-tuned lightweight model (Qwen2.5-14B).
Metrics.
Our evaluation framework considers three metrics: task success rate (TSR), execution steps (Steps), and task completion rate (TCR). TSR and Steps are the primary indicators, with TSR measuring output quality by evaluating whether the system’s response fully satisfies the user’s request (via LLM-based comparison with the ground truth), and Steps reflecting execution efficiency as the total number of LLM calls required to solve a task, counted only for successfully completed tasks. TCR serves as a supplementary metric, indicating whether the system produces a final output without prematurely terminating. Tasks are considered incomplete if they exceed the maximum allowed attempts, encounter parsing errors, or fail due to input token limits. Both TCR and TSR are reported as percentages over all evaluation tasks. All experiments details of training and inference setup provided in Appendix E.2.
4.2 Results
In this section, we present the main results on ToolBench, comparing NaviAgent with strong baselines across various model sizes and task difficulties.
Model Method Easy Medium Hard All TCR TSR Steps TCR TSR Steps TCR TSR Steps TCR TSR Steps Qwen2.5-14B ReAct 32.4 26.3 3.52 24.5 16.8 3.67 24.8 20.0 3.64 27.1 20.6 3.61 ToolLLM 56.1 30.4 4.01 53.8 19.7 4.02 38.1 11.4 4.06 51.0 21.3 4.03 $\alpha$-UMI 77.7 39.2 5.53 77.9 25.0 5.88 67.6 13.3 6.07 75.5 26.9 5.74 NaviAgent 64.2 50.7 4.32 60.6 32.7 4.32 61.9 23.8 4.47 62.0 36.4 4.34 Qwen2.5-32B ReAct 33.1 25.0 3.50 35.6 24.5 3.60 30.5 19.0 3.95 33.6 23.4 3.63 ToolLLM 40.5 31.8 3.67 48.6 30.3 3.85 49.5 23.8 4.10 46.2 29.3 3.83 $\alpha$-UMI 78.4 49.3 5.66 78.8 26.0 6.02 77.1 22.9 6.58 78.3 32.8 5.94 NaviAgent 88.5 60.8 4.33 82.2 42.8 4.70 80.0 26.7 5.11 83.7 44.9 4.60 Deepseek-V3 ReAct 46.6 36.5 3.52 58.7 38.5 3.50 48.6 23.8 3.74 52.5 34.5 3.54 ToolLLM 56.2 47.4 3.80 58.8 30.0 3.92 29.7 24.8 3.90 51.3 34.4 3.86 $\alpha$-UMI 80.8 59.7 5.95 89.4 32.9 5.95 73.0 29.5 6.64 82.9 40.7 6.06 NaviAgent 99.3 72.3 4.40 96.6 49.0 4.46 96.8 44.8 5.10 98.3 55.5 4.55
Table 1: Comparison of Baseline Frameworks on ToolBench. TCR and TSR are reported as percentages (%), and lower Steps indicates higher efficiency. The best results are marked in bold and the second-best results are marked with underline.
Figure 3: Evaluation of Frameworks on ToolBench Across Task Complexity.
Overall Performance and Efficiency.
As shown in Table 1 and Figure 3, NaviAgent consistently achieves the highest TSR across all foundation models and task complexities, with absolute values of 36.4% on Qwen2.5-14B, 44.9% on Qwen2.5-32B, and 55.5% on Deepseek-V3. Compared to the average performance of the baselines on all tasks, NaviAgent achieves substantial gains of 13.47, 16.4, and 19.0 percentage points on Qwen2.5-14B, Qwen2.5-32B, and Deepseek-V3, respectively. Meanwhile, its execution steps remain close to those of the most efficient baseline, with differences typically within one step, thereby maintaining a strong balance between solution quality and execution efficiency.
Relative Improvement and Robustness.
NaviAgent achieves an average TSR improvement of over 10 percentage points compared to $\alpha$-UMI, tthe strongest among the three baselines, across all difficulty levels, with the most significant gain of 15.33 percentage points on Deepseek-V3 for Hard tasks. We further observe that the relative drop in TSR from Easy to Hard tasks is substantially smaller for NaviAgent than for most baselines, particularly on larger foundation models. For example, on Deepseek-V3, NaviAgent’s TSR decreases by only 38.0% from Easy to Hard, while ToolLLM and $\alpha$-UMI experience drops of 47.7% and 50.6%, respectively.
Figure 4: Effect of SFT on TSR.
Adaptability through Fine-tuning.
Notably, with supervised fine-tuning, the smaller Qwen2.5-14B model achieves performance comparable to the larger 32B model (TCR 81.8% vs 83.7%, TSR 49.5% vs 44.9%, see Figure 4 and Table 2, D+N(Heur) row), indicating that fine-tuning can effectively close the gap between model sizes.
4.3 Ablation Study
To further validate the effectiveness of each component in our framework, we conduct two sets of ablation studies.
Multi-Path Decider Analysis.
We analyze the execution paths taken by only Decider in successful cases on ToolBench, focusing on the Deepseek-V3 model. Specifically, we categorize the proportion of cases resolved via Clarification (intent clarification to seek additional details from the user), Re-retrieval (recovering from initial tool retrieval failures by invoking alternative APIs), and Normal (tasks completed successfully in a single attempt without clarification or re-retrieval). Results are summarized in Figure 5, demonstrating that multi-path planning enables robust error recovery and flexible intent handling, contributing to overall performance gains.
Figure 5: Comparison of TSR Distribution Between Multi-Path Decider and Baselines.
Graph-Encoded Navigator Analysis.
Table 2 indicates that introducing the Navigator module consistently improves TSR across all tasks and model sizes, with our heuristic search strategy providing additional gains, achieving an average TSR improvement of 2.4 points over Alpha-Beta search. Notably, these benefits become more pronounced as model capacity increases. On hard tasks, the addition of the Navigator yields relative TSR improvements of 27.3% for Deepseek-V3 and 36.2% for GPT-4o. This demonstrates that graph-based planning and efficient search more effectively unlock the reasoning and compositional capabilities of larger models, especially for complex toolchain tasks. Consistent results are also observed on API-Bank (see Table 5). Additional statistics on TDHG structure and link prediction are provided in Table 6.
Model Method Easy Medium Hard All TCR TSR Steps TCR TSR Steps TCR TSR Steps TCR TSR Steps Qwen2.5-14B Decider 54.7 41.2 4.35 58.2 29.3 4.30 66.7 21.0 4.55 59.0 31.2 4.36 D+N(Alpha) 58.8 48.0 4.21 61.5 31.7 4.33 53.3 16.2 4.28 58.8 33.4 4.27 D+N(Heur) 64.2 50.7 4.32 60.6 32.7 4.32 61.9 23.8 4.47 62.0 36.4 4.34 Qwen2.5-32B Decider 84.5 52.0 4.32 81.3 40.9 4.83 90.5 24.8 5.70 84.4 40.8 4.74 D+N(Alpha) 83.1 51.4 4.31 85.1 41.3 4.87 80.0 31.4 5.04 83.3 42.3 4.68 D+N(Heur) 88.5 60.8 4.33 82.2 42.8 4.70 80.0 26.7 5.11 83.7 44.9 4.60 Deepseek-R1-32B Decider 94.6 40.5 5.01 92.8 34.1 5.49 93.3 27.6 5.29 93.5 34.7 5.27 D+N(Alpha) 92.6 51.4 4.96 93.3 38.0 5.17 91.4 21.9 5.60 92.6 38.6 5.13 D+N(Heur) 93.2 51.4 4.98 92.3 38.0 5.33 88.6 33.3 5.27 91.8 41.2 5.18 Deepseek-V3 Decider 99.9 71.6 4.26 99.0 47.1 4.88 99.9 35.2 4.95 99.6 52.3 4.62 D+N(Alpha) 93.2 71.6 4.26 95.7 50.5 4.52 93.3 33.3 4.64 94.4 53.4 4.42 D+N(Heur) 99.3 72.3 4.40 96.6 49.0 4.46 96.8 44.8 5.10 98.3 55.5 4.55 GPT-4o Decider 99.8 67.6 4.05 98.6 43.9 4.60 99.0 35.9 4.95 99.1 49.7 4.42 D+N(Alpha) 99.9 76.4 4.08 99.5 45.3 5.24 98.1 41.4 5.59 99.3 54.4 4.78 D+N(Heur) 99.9 72.1 4.05 95.7 50.0 4.78 99.0 48.9 4.96 97.8 56.8 4.52 Qwen2.5-14B(SFT) Decider 77.7 55.4 4.80 78.8 46.2 4.78 75.2 30.5 5.57 77.7 45.6 4.91 D+N(Alpha) 85.8 64.9 4.48 78.4 39.9 4.59 78.1 39.0 5.26 80.7 47.7 4.67 D+N(Heur) 83.8 64.2 4.68 81.7 49.0 4.73 79.0 29.5 5.66 81.8 49.5 4.84
Table 2: Impact of Navigator and Search Algorithms on ToolBench. Decider refers to using only the decider module. D+N(Alpha) denotes Decider + Navigator with Alpha-Beta search. D+N (Heur) denotes Decider + Navigator with our proposed heuristic search. Metrics are reported as in Table 1.
5 Conclusion
We present NaviAgent, a bilevel planning framework for compositional tool reasoning that enhances large language models’ planning and execution abilities. By combining a decider and a graph-based navigator with heuristic search, NaviAgent delivers robust improvements in task completion and success rates across diverse tasks and model sizes, while maintaining competitive efficiency. Experiments on ToolBench and API-Bank demonstrate that NaviAgent is particularly effective for complex, multi-tool tasks and larger models, consistently achieving strong TSR gains.
Limitations.
Despite its strengths, NaviAgent has several limitations. First, integration challenges from heterogeneous tool interfaces remain to be improved, which could be mitigated by adopting unified protocols like MCP. Second, the single-agent design limits collaboration and generalization in broader or multimodal tasks. Finally, although NaviAgent incorporates adaptive graph evolution, highly dynamic or noisy environments could still challenge robust graph construction and heuristic search. Addressing these issues will be important for future work.
Appendix A Graph Construction
Original API Original Parameter Parameter Description Standardized Parameter Cluster ID get_locations name Name of the city. city_name 1 get_hospital_list city The city where the hospital is located. city_name 1 get_hospital_list name Name of the hospital. hospital_name 2 find_cheapest_prescription city_name The name of the city where the user wants to search for the medication. city_name 1
Table 3: Standardization of API Parameter
Appendix B Details of Graph Method
B.1 HGT Network
This section provides a detailed description of the feature construction, network architecture, and link prediction head used in our heterogeneous graph transformer (HGT) for tool dependency modeling, supplementing the main text.
Feature Fusion.
Each node $v$ is initialized by its semantic and structural features:
$\mathbf{h}_{v}={BGE}(x_{v})\oplus\sigma(n_{v}^{succ})\oplus\sigma(r_{v}^{succ})\oplus\sigma(deg_{v}^{in})\oplus\sigma(deg_{v}^{out})$ (11)
where ${BGE}(x_{v})$ encodes the node description $d_{v}$ using BGE-Large-en-V1.5, $n_{v}^{succ}$ and $n_{v}^{fail}$ are the counts of successful and failed invocations for node $v$ (computed from historical invocation logs), $r_{v}^{succ}=n_{v}^{succ}/(n_{v}^{succ}+n_{v}^{fail})$ denotes the successful ratio, and $deg_{v}^{in}$ and $deg_{v}^{out}$ are the in-degree and out-degree of node $v$, respectively.
Node Encoder.
To project heterogeneous nodes into a unified embedding space, we apply type-specific linear transformations, followed by non-linear activation and normalization:
$\mathbf{h}_{v}^{\prime}={LayerNorm}\left({LeakyReLU}\left(\mathbf{W}_{\tau(v)}\mathbf{h}_{v}+\mathbf{b}_{\tau(v)}\right)\right)$ (12)
where $\mathbf{W}_{\tau(v)}$ and $\mathbf{b}_{\tau(v)}$ are the learnable weight matrix and bias for node type $\tau(v)\in\{api,param\}$, respectively.
WeightedHGTConv Layer.
We stack two multi-head heterogeneous graph transformer (HGT) layers (each with 8 attention heads) to aggregate information from the 2-hop neighborhood. For a center node $v$ and its neighbor $u\in N_{r}(v)$ under edge type $r$, the attention coefficient at head $k$ is computed as:
$\alpha_{uv}^{(k,r)}={softmax}_{u\in\mathcal{N}_{r}(v)}\left(\frac{(\mathbf{W}_{Q}^{(k,r)}\mathbf{h}_{u}^{\prime})^{\top}(\mathbf{W}_{K}^{(k,r)}\mathbf{h}_{v}^{\prime})}{\sqrt{d_{k}}}+\mathbf{b}_{r}^{(k)}+\tilde{w}_{uv}\right)$ (13)
where $\mathbf{W}_{Q}^{(k,r)}$ and $\mathbf{W}_{K}^{(k,r)}$ are the query and key projection matrices for head $k$ and relation $r$, $\mathbf{b}_{r}^{(k)}$ is an edge-type-specific bias, $\tilde{w}_{uv}$ is the statistical edge weight from node $u$ to $v$ (see Eq. 4, where $\tilde{w}_{ij}$ is defined for nodes $v_{i}$ and $v_{j}$), and $d_{k}=d/8$ is the dimension per head. The normalization ${softmax}_{u\in\mathcal{N}_{r}(v)}$ is performed over all neighbors $u$ of $v$ under relation $r$. The output embedding for node $v$:
$\mathbf{h}_{v}^{\prime\prime}={LayerNorm}\left(\mathbf{h}_{v}^{\prime}+{LeakyReLU}\left(\mathbf{W}_{o}\cdot{Concat}\left[\sum_{r\in R}\sum_{u\in\mathcal{N}_{r}(v)}\alpha_{uv}^{(k,r)}\mathbf{W}_{V}^{(k,r)}\mathbf{h}_{u}{\prime}\right]_{k=1}^{8}\right)\right)$ (14)
where $\mathbf{W}_{V}^{(k,r)}$ is the value projection for head $k$ and relation $r$, $\mathbf{W}_{o}\in\mathbb{R}^{8d_{k}\times d}$ is the output projection, and $Concat[\cdot]_{k=1}^{8}$ denotes concatenation of outputs from all heads.
Link Prediction.
Given the final node embeddings, the link probability between node $u$ and node $v$ is computed as:
$p_{uv}=\sigma\left(\mathbf{W}_{p}\cdot{Concat}(\mathbf{h}_{u}^{\prime\prime},\mathbf{h}_{v}^{\prime\prime})+\mathbf{b}\right)$ (15)
where $\mathbf{W}_{p}$ and $\mathbf{b}$ are learnable parameters, and $\sigma(\cdot)$ denotes the sigmoid function.
This completes the detailed description of our HGT-based network architecture.
B.2 Training Objective
This section provides the full definitions of the loss functions and training strategies used for graph optimization.
Cross-Entropy Loss.
The cross-entropy loss encourages the model to match predicted link probabilities with observed edge weights:
$\mathcal{L}_{\mathrm{CE}}=-\frac{1}{|\mathcal{E}|}\sum_{(u,v)\in\mathcal{E}}\left[\tilde{w}_{uv}\log p_{uv}+(1-\tilde{w}_{uv})\log(1-p_{uv})\right]$ (16)
where $p_{uv}$ is the predicted link probability, $\tilde{w}_{uv}$ is the statistical edge weight serving as a soft label, and $\mathcal{E}$ denotes the set of all edges in the graph.
Adaptive Margin Loss.
To further enhance representation discrimination, we introduce an adaptive margin loss that dynamically assigns larger separation intervals to node pairs with higher edge weights (i.e., $\tilde{w}_{uv}\rightarrow 1$), encouraging the model to focus on learning critical dependencies. For each positive edge $(u,v)^{+}\in\mathcal{E}^{+}$, $k$ negative edges $\{(u_{j},v)\}_{j=1}^{k}$ are sampled to construct positive and negative pairs for the margin loss.
$m_{uv}=m_{0}\left(1+\sigma(\tilde{w}_{uv})\right)$ (17)
$\mathcal{L}_{\mathrm{Margin}}=\frac{1}{|\mathcal{E}^{+}|}\sum_{(u,v)^{+}\in\mathcal{E}^{+}}\frac{1}{k}\sum_{j=1}^{k}\left[m_{uv}-s(u,v)^{+}+s(u_{j},v)^{-}\right]_{+}$ (18)
where $m_{0}$ is a base margin, $\sigma(\cdot)$ denotes the sigmoid function, $\tilde{w}_{uv}$ is the statistical edge weight, $s(u,v)$ measures the embedding similarity, $(u,v)^{+}$ represents a positive edge, $(u_{j},v)^{-}$ denotes a negative sample, and $[\cdot]_{+}$ is the hinge function, and $\mathcal{E}^{+}$ is the set of positive edges.
Hybrid Loss.
The final training objective is a weighted sum of the two losses:
$\mu_{t}=\mu_{0}\cdot\gamma^{t},\ \gamma\in(0,1)$ (19)
$\mathcal{L}=\mu_{t}\cdot\mathcal{L}_{CE}+(1-\mu_{t})\cdot\mathcal{L}_{\mathrm{Margin}}$ (20)
where $\mu_{t}$ is the weight for the cross-entropy loss at epoch $t$, $\mu_{0}$ is the initial weight, and $\gamma$ is a decay factor controlling the rate at which the contribution of the cross-entropy loss decreases over training.
B.3 Graph Search Algorithm
This section provides detailed descriptions of the Alpha-Beta pruning and hybrid heuristic search algorithms, including all parameter settings, dynamic thresholding strategies, and algorithmic pseudocode.
Alpha-Beta Pruning.
We adapt the Alpha-Beta pruning algorithm for backward search over the tool dependency graph $\mathcal{G}=(V,E,W)$, parameterized by a quintuple $(\alpha,\beta,\mathcal{H},\mathcal{D},\mathcal{C})$, where $\alpha\in\mathbb{R}^{+}$ (initialized as $\alpha_{0}=0.4$) is the lower-bound threshold for acceptable path scores, and $\beta\in\mathbb{R}^{+}$ (with $\beta_{0}=0.9$) is the upper-bound for candidate evaluation. The dynamic threshold function $\mathcal{H}(d)=\max(0.3,0.5\times 0.9^{d})$ applies exponential decay to balance search depth $d$ and semantic relevance. The depth attenuation factor $\mathcal{D}(d)=1/(1+\sqrt{d})$ penalizes longer paths. The connectivity constraint $\mathcal{C}(u,v_{t})=\mathrm{PathLength}(u,v_{t})\leq 5$ ensures that generated subgraphs remain compact, where $v_{t}$ denotes the target node (either an API node or a parameter node). The parametric scoring function is defined as:
$S_{uv}=\frac{w_{uv}+\mathbb{I}(u\rightarrow v_{t}^{\mathrm{api}})w_{u\rightarrow v_{t}^{\mathrm{api}}}+\mathbb{I}(u\rightarrow v_{t}^{\mathrm{param}})w_{u\rightarrow v_{t}^{\mathrm{param}}}}{3}\times\mathcal{D}(d)$ (21)
where $w_{uv}$ is the direct edge weight from node $u$ to its predecessor $v$ (see Section 3.2.2), $w_{u\rightarrow v_{t}^{\mathrm{api}}}$ and $w_{u\rightarrow v_{t}^{\mathrm{param}}}$ denote the edge weights from $u$ to the target API node $v_{t}^{\mathrm{api}}$ and target parameter node $v_{t}^{\mathrm{param}}$, respectively, included only if the corresponding indicator function $\mathbb{I}(\cdot)$ is active.
During reverse depth-first search, we apply two pruning rules: Alpha-pruning is triggered at parameter nodes when $S_{uv}<\mathcal{H}(d)$ and $S_{uv}<\alpha$, while Beta-pruning is triggered at API nodes when $S_{uv}>\beta$. To further improve efficiency, the pruning thresholds are dynamically adjusted via $\alpha^{\prime}=\max(\alpha,S_{uv}\times 0.85)$ and $\beta^{\prime}=\min(\beta,S_{uv}\times 1.15)$, reducing the search time complexity from $O(b^{k})$ to $O((\sqrt{b})^{k})$ knuth1975analysis , where $b$ is the branching factor and $k$ is the maximum search depth. See Algorithm 1 for details.
Algorithm 1 Alpha-Beta Backward Pruning
Input G, $v_{target}$, $\alpha_{init}=0.4$, $\beta_{init}=0.9$, $d_{max}=5$
Output $G_{sub}$
Initialize queue $Q$ with $v_{target}$, $V_{visited}$ = {${v_{t}}$}
while Q not empty do
     v = Q.pop()
     for $p\in predecessors(v)$ do
         if $p$ not in $V_{visited}$ then
              $s=Score(p\rightarrow u,p\rightarrow v_{target},p\rightarrow v_{target\_param})$
              d = current_depth(p)
              $\mathcal{H}(d)=\max(0.3,0.5\times 0.9^{d})$
              if $p\in V_{param}$ then
                  if $s<\mathcal{H}(d)\land s<\alpha$ then
                       continue
                  end if
                  if $s>\beta$ then
                       break
                  end if
              end if
              $\alpha=max(\alpha,s\times 0.85)$
              $\beta=max(\beta,s\times 1.15)$
              $V_{visited}.add(p)$
              $Q.append(p)$
         end if
     end for
end while
$V_{sub}=\{{v|v\in V_{visited}\land{PathLength}(u,v_{target})\leq 5}\}$
return $G_{sub}=(V_{sub},E)$
Figure 6: Alpha-Beta Backward Pruning
Heuristic Graph Search with Dynamic Pruning.
The algorithm is parameterized by a sextuple $(\mathcal{T}_{0},\eta,\mathcal{P},d_{\max},\mathcal{M}_{\theta},\mathcal{F}_{\omega})$ (see Algorithm 2), where $\mathcal{T}_{0}=300$ is the initial temperature that determines the probability of accepting suboptimal solutions and balances exploration and exploitation, $\eta=0.7$ is the cooling rate that controls the annealing schedule $\mathcal{T}_{k+1}=\eta^{1+k/5}\mathcal{T}_{k}$, $\mathcal{P}=40$ is the population size, $d_{\max}=4$ is the maximum search depth, and $\mathcal{M}_{\theta}$ is a temperature-sensitive mutation operator with adaptive intensity $\theta=\lfloor\mathcal{T}/100\rfloor$. Candidate solutions are evaluated using a composite fitness function:
$\mathcal{F}_{\omega}=0.35\mathcal{C}_{c}+0.15\log(1+\rho_{p})+0.3\mathcal{D}_{c}+0.15\mathcal{W}_{n}+0.05\mathcal{C}_{p}$ (22)
where $\mathcal{C}_{c}$ (node compactness) measures the closeness centrality of API nodes, $\rho_{p}$ (parameter density) is the ratio of parameter nodes within the subgraph to promote concise yet informative solutions, $\mathcal{D}_{c}=0.2e^{-d/10}+0.8e^{-n/8}$ (depth penalty) penalizes overly deep or complex dependency structures, with $d$ as the average depth and $n$ as the total node count, $\mathcal{W}_{n}$ (weight quantification) encourages solutions with higher cumulative edge weights, and $\mathcal{C}_{p}$ (path complexity) evaluates structural simplicity, favoring solutions with less intricate connectivity.
Algorithm 2 Hybrid Heuristic Pruning Algorithm
Dependency graph $G$, target API set $A$, initial temperature $\mathcal{T}_{0}=300$, cooling rate $\eta=0.7$, population size $\mathcal{P}=40$, maximum search depth $d_{max}=4$
2:Optimized dependency subgraph $G^{*}$
Initialize optimized subgraph set $S\leftarrow\emptyset$
4:for each target API $a\in A$ do
     Set temperature $\mathcal{T}\leftarrow\mathcal{T}_{0}$
6:     Generate initial population $Pop$ (size $\mathcal{P}$) for API $a$
     Set iteration count $k\leftarrow 0$
8:     while $\mathcal{T}>1$ and $k\leq 10$ do
         Evaluate fitness $\mathcal{F}_{\omega}$ for each chromosome in $Pop$
10:         Select elite chromosomes (top 60% based on fitness)
         Generate offspring via crossover operation
12:         Apply temperature-sensitive mutation $\mathcal{M}_{\theta}$ with intensity $\theta=\lfloor\mathcal{T}/100\rfloor$
         Update population $Pop$ with offspring
14:         Update temperature: $\mathcal{T}\leftarrow\eta^{1+k/5}\mathcal{T}$
         Increment iteration count $k\leftarrow k+1$
16:     end while
     Select best chromosome from $Pop$ based on $\mathcal{F}_{\omega}$
18:     Decode chromosome and build API-specific subgraph $G_{a}$
     Add subgraph $G_{a}$ to optimized subgraph set $S$
20:end for
Merge all subgraphs in $S$ into final optimized subgraph $G^{*}$
22:return $G^{*}$
Figure 7: Hybrid Heuristic Pruning Algorithm
Appendix C Cases
The following three cases exemplify the Decider-Navigator collaborative mechanism through four core actions executed by the Decider: 1) Direct Response: resolves user queries using pre-trained knowledge. 2) Intent Clarification: initiates interactive dialogue to disambiguate vague requests. 3) Tool Retrieval: collaborates with the Navigator module to generate a pruned tool dependency subgraph, which is returned as executable toolchains. 4) Tool Call: using above dependency subgraph, Executes API with parameter validation and state monitoring. This design achieves centralized decision control through the Decider’s orchestration authority while enabling dynamic resource optimization via the Navigator’s graph-based toolchain generation, ensuring both operational efficiency and system integrity in complex task environments.
C.1 Case 1
Query
  • Could you provide me with information about gastroenteritis? Additionally, please help me log my health data from March 5, 2025.
First Round: Intent Analysis
  1. 1.
    Gastroenteritis Inquiry
    ACTION: Direct Response
    CONTEXT:
    "analysis": "1. Common sense judgment: Gastroenteritis is a medical
    condition within general knowledge. 2. No API required for definition
    explanation. 3. Parameter verification not applicable",
    "action": "direct_answer",
    "intent": "1. Provide medical definition of gastroenteritis 2. User’s
    explicit request for condition explanation",
    "target_api": "",
    "params": {},
    "recall_description": "",
    "answer": "Gastroenteritis is inflammation of the stomach and intestines,
    typically resulting from bacterial toxins or viral infection. Common
    symptoms include diarrhea, vomiting, and abdominal cramps."
  2. 2.
    Health Data Logging
    ACTION: Tool Retrieval
    CONTEXT:
    "analysis": "1. Health data recording requires specialized API 2. No
    matching API in candidate list 3. Critical parameters provided:
    user_id=12345, blood_pressure=130/80, heart_rate=90, timestamp=2025-03-05",
    "action": "retrieve_api",
    "intent": "1. Store user health metrics 2. Provided parameters: blood
    pressure, heart rate, user ID, timestamp",
    "target_api": "",
    "params": {},
    "recall_description": "HealthDataAPI(description: Records user health
    metrics, input: user_id:string/unique identifier,
    blood_pressure:string/systolic & diastolic, heart_rate:integer/bpm,
    timestamp:datetime/measurement time; output: record_id:string/new entry
    identifier)",
    "answer": "I need to retrieve a health data recording API to complete this
    action. Should I proceed with API retrieval?"
    Then, the BGE model retrieves three candidate APIs (api-RecordHealthData, api-QueryHealthData, api-QueryRegistration) based on the recall description above, after which the Navigator module is invoked to acquire the corresponding call chains for these APIs (see Figure 8).
    Figure 8: Pruned Tool Dependency Subgraph of Case1
Second Round: Iterative API Chain Execution
In this round, the Decider module selects and executes the optimal API chain, with api-RecordHealthData designated as the target API.
ACTION: Tool Call
CONTEXT:
  "action": "call_api",
  ’api_name’: ’api-QueryHealthData’
  "params": {
    "user_id": "12345",
    "start_time": "2025-03-05",
    "end_time": "2025-03-05"
TOOL RESPONSE:
{{name:blood_pressure,value:130/80}
{name:heart_rate,value:90}}
Third Round: Iterative API Chain Execution
ACTION: Tool Call
CONTEXT:
  "action": "call_api",
  ’api_name’: ’api-RecordHealthData’
  "params": {
    "user_id": "12345",
    "time": "2025-03-05 14:30:00",
    "health_data": "[{’name’:’blood_pressure’,’value’:’130/80’},
                    {’name’:’heart_rate’,’value’:’90’}]"
TOOL RESPONSE:
Finished
Fourth Round: API Output Verbalization
ACTION: Direct Response
CONTEXT:
  "analysis": "1. API parameters verified and complete
               2. Successful data storage confirmation",
  "action": "direct_answer",
  "answer": "Your health data has been successfully recorded..."
Last Round: Response Synthesis
ACTION: Direct Response
CONTEXT:
1.Inflammation of gastrointestinal tract
    Common causes: Viral/bacterial infections
    Key symptoms: Diarrhea, vomiting, abdominal cramps
    Management: Hydration and rest
2. Health Data Record Confirmation:
    All parameters validated successfully through api-RecordHealthData. No data
    integrity issues detected.
C.2 Case 2
Query
  • Is Beijing in China? What is the current weather forecast for the location at latitude 40.7128 and longitude -74.0060 in metric units?
First Round: Intent Analysis
  1. 1.
    Common Sense
    ACTION: Direct Response
    CONTEXT:
    {’dependency_rank’: 1,
      ’intent’: ’Is Beijing in China?’,
      ’analysis’: ’1. The specific problem addressed by this action is to determine
      if Beijing is located within China. 2. The user input directly asks about the
      geographical location of Beijing. 3. This is a common sense question and does
      not require an API call. 4. No parameters are needed for this decision.’,
      ’action’: ’direct_answer’,
      ’answer’: ’Yes, Beijing is in China.’}
  2. 2.
    Get Weather
    ACTION: Tool Retrieval
    CONTEXT:
    {’dependency_rank’: 2,
    ’intent’: ’What is the current weather forecast for the location at latitude
    40.7128 and longitude -74.0060 in metric units?’,
    ’analysis’: ’1. The specific problem addressed by this action is to retrieve
    the current weather forecast for a specific location using latitude and
    longitude coordinates. 2. This is not a common sense question and requires an
    external API call. 3. The API matching phase will prioritize APIs that can
    provide weather forecasts based on geographic coordinates. 4. Required
    parameters are latitude and longitude, which are provided in the user input.’,
    ’action’: ’retrieve_api’,
    ’recall_description’: ’WeatherAPI(description: Retrieve current weather
    conditions by geographic coordinates, input: latitude:float/latitude
    coordinate; longitude:float/longitude coordinate; output:
    temperature:float/current temperature; humidity:float/current humidity;
    wind_speed:float/current wind speed)’}
    Then, the BGE model retrieves three candidate APIs (api-current_weather_data_of_a_loca-
    tion_for_weather, api-by_geographic_coordinates_for_rapidweather, api-current_for_foreca
    _weather) based on the recall description above, after which the Navigator module is invoked to acquire the corresponding call chains for these APIs (see Figure 9).
    Figure 9: Pruned Tool Dependency Subgraph of Case2
Second Round: Iterative API Chain Execution
In this round, the Decider module selects and executes the optimal API chain, with api-by_geographic
_coordinates_for_rapidweather designated as the target API.
ACTION: Tool Call
CONTEXT:
  "action": "call_api",
  ’api_name’: ’api-by_geographic_coordinates_for_rapidweather’
  ’params’: {’param-lat’: ’40.7128’, ’param-lon’: ’-74.0060’}
TOOL RESPONSE:
{temperature:11.0C, feels:8.0C, wind_speed:5.1 m/s}
Third Round: API Output Verbalization
ACTION: Direct Response
CONTEXT:
  ’analysis’: "1. The specific problem addressed by this action is to retrieve
  the current weather forecast for a specific location using the provided
  latitude and longitude. 2. The user input directly provides the necessary
  parameters (latitude and longitude) and specifies the need for metric units.
  3. The target API ’api-current_weather_data_of_a_location_for_weather’ has
  already been successfully called and its result is available. 4. The required
  parameters (lat and lon) are complete and the API call has been confirmed
  successful.",
  ’action’: ’direct_answer’,
  ’answer’: ’The current weather in New York is clear sky with a temperature of
  11.0C, feels like 8.0C, and the wind speed is 5.1 m/s from the west-
  southwest.’}
Last Round: Response Synthesis
ACTION: Direct Response
CONTEXT:
Beijing is in China. The current weather in New York is clear sky with a
temperature of 11.0C, feels ike 8.0C, and the wind speed is 5.1 m/s from the west-
southwest.
C.3 Case 3
Query
  • Could you tell me what the weather will be like tomorrow?
First Round: Intent Analysis
  1. 1.
    Intent Clarification
    ACTION: Intent Clarification
    CONTEXT:
    "analysis": "1. Weather data needs to be queried in real time -> not common
[…]
    "action": "clarify_intent",
    "recall_description": "",
    "answer": "Which city do you want to query tomorrow’s weather?"
  2. 2.
    User Answer
    User Answer:
        I’m in BeiJing.
Second Round: Intent Analysis
  1. 1.
    Get Weather
    ACTION: Tool Retrieval
    CONTEXT:
    {’dependency_rank’: 1,
      ’intent’: ’1. Query the weather forecast for tomorrow in Beijing 2. Extract
      location: Beijing and time: tomorrow from user input’,
      ’analysis’: ’1. The specific problem addressed by this action is to retrieve
      the weather forecast for tomorrow in Beijing. 2. The user input directly
      provides the location (Beijing) and the time frame (tomorrow). 3. Since no
      candidate APIs are available, the system needs to retrieve an appropriate API
      for weather forecasting. 4. There are no parameters provided by the user that
      can be used directly with an API, so the system must retrieve an API that can
      accept location and time as parameters.’,
      ’action’: ’retrieve_api’,
      ’recall_description’: ’WeatherForecastAPI(description: Retrieve weather
      forecast for a given location and date, input: location:string/city name;
      date:date/forecast date; output: weather:string/weather condition,
      temperature:float/forecast temperature)’}
    Then, the BGE model retrieves three candidate APIs (api-getweatherforecast_for_apjoy_
    weather_forecast, api-weather_report_for_the_weather_api, api-location_info_for_foreca
    _weather) based on the recall description above, after which the Navigator module is invoked to acquire the corresponding call chains for these APIs (see Figure 10).
    Figure 10: Pruned Tool Dependency Subgraph of Case3
Third Round: Iterative API Chain Execution
In this round, the Decider module selects and executes the optimal API chain, with api-getweatherforecast_for_apjoy_weather_forecast designated as the target API.
ACTION: Tool Call
CONTEXT:
  "action": "call_api",
  ’api_name’: ’api-getweatherforecast_for_apjoy_weather_forecast’
  ’params’: {’location’: ’Beijing’}
TOOL RESPONSE:
{temperature:25C, humidity:60%}
Last Round: API Output Verbalization
ACTION: Direct Response
CONTEXT:
  The weather forecast for tomorrow in Beijing is partly cloudy with a
  temperature of 25C and humidity of 60%.
Appendix D Data Generation
Problem and API Dependency Chain Generation Prompts: Automatically construct a chain of calls with strict parameter matching based on a randomful tree structure. This process involves tracing initial parameters from the leaf nodes and validating results at the root node to ensure that the generated problem aligns with the authentic API dependency logic of real-world scenarios.
1. Core Requirements:
   - Generate a natural-language question where:
      Must explicitly contain initial parameters for leaf-node APIs
      Implicitly requires chained API calls from leaf to root node
      Root node API’s output directly resolves the user’s problem
2. Dependency Chain Rules:
   - Build parameter-passing paths where:
      Parent API outputs must exactly match child API inputs (same
       parameter names & data types)
      Root node API must be called last in the chain
      The output of every leaf-node API must be utilized in downstream
       APIs or final results.
      All input values must originate from either:
        Explicitly stated in the question context
        Generated by previous API outputs (no synthetic values)
3. Parameter Constraints:
   - Enforce strict value inheritance:
      Path/query parameters must use verbatim values from:
       - User’s question text
       - Preceding API response.data fields
      Prohibit value transformation/format conversion
   - Root API output must contain realistic values matching its schema
4. Validation Requirements:
   - Reject generation if:
      Missing parameter dependency between APIs
      Input sources can’t be traced to question/prior responses
      Output fields don’t fulfill next API’s input requirements
5. Response Structure:
  "query": "<Real-world scenario requiring sequential API calls>",
  "answer": "<Solution derived from root API output>",
  "call_chains": [
      "api_name": "<Leaf-node API>",
      "input": {
        "<param>": "<value explicitly stated in user query or previous
                   API output>"
      "output": {
        "status": "success",
        "data": {"<field>": "<output used by next API>"}
      "api_name": "<Root-node API>",
      "input": {
        "<param>": "<value from previous API output>"
      "output": {
        "status": "success",
        "data": {"<field>": "<realistic resolution to query>"}
The API dependency tree structure is as follows:
Appendix E Implementation Details
E.1 Dataset
This table 4 system displays the sample counts of the ToolBench and API-Bank datasets, as well as the distribution of their difficulty levels.
Dataset Easy Medium Hard Total
API-Bank 57 176 211 444
ToolBench 148 208 105 461
Table 4: Dataset Samples and Difficulty Distribution
E.2 Training
We fine-tunetajbakhsh2016convolutional our model using Qwen2.5-14B model with full parameter tuning. The model is trained with a maximum sequence length of 8192. We utilize a learning rate of 2e-5 and employ the AdamW optimizer with a cosine learning rate scheduler. The training process includes 10 epochs with a per-device batch size of 1 for both training and evaluation. Gradient checkpointing is enabled to reduce memory usage, and gradient accumulation is set to 4 steps to effectively manage smaller batch sizes. We apply a weight decay of 0.01 and set the maximum gradient norm to 1 for stable training. A warmup ratio of 0.03 is used to gradually increase the learning rate at the beginning of training. The training is executed on 8 Ascend 910B 64G GPUs within 10 hours. The DeepSpeedrasley2020deepspeed library is leveraged for efficient distributed training.
E.3 Inference
E.3.1 NaviAgent Inference Prompts
Inference prompts are based on intent decomposition and dependency prioritization to achieve automatic parameter completion and error handling. They generate standardized JSON responses through hierarchical decision-making.
You are an intelligent API coordination system. Respond strictly according
to the following rules:
# Decision Architecture
1. **Intent Analysis**
   - Decompose compound requests into independent ordered sub-intents
      Sequential dependencies first, Must execute in declared order
      Parallelizable sub-intents last
      Dependency_rank numbering for ordered execution
   - Validate parallel execution eligibility:
      No overlapping data requirements
      No sequential dependencies
      Distinct parameter sets
2. **Atomic Action Formation**
     For each validated sub-intent:
       - Create self-contained decision unit, action must implement full
       Decision Logic Flow
       - Maintain state separation between parallel processes
       - Focus analysis scope per sub-intent
       - Each action’s analysis focuses only on its own intent
       - Each action analysis only solves one intent
       - Must execute each action in declared order
# Decision Logic Flow
1. **Common Sense Judgment Phase**
   - Input question -> Knowledge base matching
    Belongs to common sense -> action=direct_answer
    Requires external data -> Proceed to Phase 2
2. **API Matching Phase**
   1. If candidate_apis is empty -> action=retrieve_api
   2. Match intent with API list:
     API prioritization:
          - Complete parameters from user input
          - Minimal missing parameters
          - Shortest dependency chain
     API matching success:
        - Validate Observation in user input to confirm target API success:
           -> If successful -> action=direct_answer
           -> No explicit success indication:
            a) Complete parameters -> action=call_api (execute based on 3.1
            dependency resolution)
                - If Rule 3.1c applies -> action=direct_answer
            b) Missing parameters -> Proceed to Phase 3
     API matching failed -> action=retrieve_api
3. **Parameter Completion Phase**
   - Check required parameter set:
     All parameters ready -> action=call_api
     The target API does not require parameters -> action=call_api
     Missing parameters exist:
       a) Can be completed via dependent APIs -> Execute Rule 3.1
       b) Use Retrieval APIs resolve parameter deficiencies in API
          dependencies -> action=retrieve_api
       c) Requires human confirmation -> action=clarify_intent
# Technical Rule
## 3.1 Dependency Resolution Rules
    a) Check required parameters of target API, first call dependent APIs.
    b) For each missing parameter, select APIs from dependencies not marked
       as failed.
    c) If an input parameter of an API is unavailable, use retrieve_api to
       call another API that generates it from known parameters.
       -> action=retrieve_api
    d) Success propagation: Completed dependency chain
       -> action=direct_answer
## 3.2 Known Failure Handling
    a) Failed APIs are recorded in failed_apis
    b) Prioritize non-failed candidate APIs
# Response Specification (Mandatory JSON Format)
  "dependency_rank": 1,
  "intent": "1. <precisely describe the specific problem addressed by the
        current action>
             2. <extract data segments directly related to the subtask from
        user input>",
  "analysis":
        "<Four-level reasoning:
        1.Explicitly state the specific decision-making sub-intent
          addressed by this action
        2.Common sense judgment basis
        3.API matching logic (if applicable)
        4.Parameter completeness verification>",
  "action": "call_api|direct_answer|retrieve_api|clarify_intent",
  "target_api": "API name (mandatory for call_api)",
  "params": {"parameter": "value (mandatory for call_api)"},
  "recall_description":
        "When action=retrieve_api: Use ’APIName(description: API
        functionality, input: param:type/description; output:
        param:type/description)’ format with only core parameters (e.g.,
        ’StockAPI(description: Query stock price by symbol, input:
        symbol:string/stock symbol; output: price:float/current price)’)",
  "answer": "When actionin[direct_answer,clarify_intent]: Natural language
        response (interrogative sentences required)"
# Response Specification
Added constraint:
- JSON array items MUST be sorted by dependency_rank in ascending order
- Sibling sub-intents should have consecutive ranks
# Termination Conditions
[OR]Generate final answer
[OR]Target API must be executed successfully, as shown in the status
# Enforcement Constraints
1. Parameter names must strictly match API documentation
2. The ’answer’ field for clarify_intent must contain question words
3. Prioritize calling parent node APIs
4. When action in [retrieve_api]:
    - The recall_description field serves exclusively as an API retrieval
      identifier from predefined repositories.
    - parameter descriptions must distinguish between input and output
      parameters, retaining only essential parameters
    - Each recall_description can only recall one api,multiple APIs require
      multiple actions.
5. APIs absent from Candidate APIs MUST NOT be invented
6. When action=call_api is permitted only when candidate APIs exist and the
   target_api is present in the candidate APIs.
7. The "action" field must be strictly limited to one of the following four
   predefined operation types: call_api, direct_answer,retrieve_api or
   clarify_intent.
8. Use retrieve_api only when:
    - Required parameters unavailable in call_api action
9. Use call_api only when:
    - The target_api is not in the list of successfully executed APIs
# Candidate API Information:
E.3.2 Input Generation Prompts
Input generation prompts: Integrate current queries with observational data to formulate the final input, ensuring informational completeness.
[…]
following data:
{observation} :
    Requirements:
    1. Integrate all available data
    2. Indicate data limitations (if any failed APIs exist)
    3. Use natural and fluent English
E.3.3 API Simulator Prompts
API simulator prompts are based on historical data reuse (Case1) and intelligent simulation generation (Case2/3). They achieve automated emulation of API chains through standardized JSON responses. The priority strategy is as follows: historical matching > structural cloning > contextual simulation.
Act as an API Chain Simulator to generate responses based on historical call
chains.
Follow these rules strictly:
Operation Rules:
1. Request Processing Logic
   - CASE 1: Existing API + Identical Inputs
      Return historical outputs verbatim
      Set {"status": "success", "type": "success"}
   - CASE 2: New API
      Create mock data matching input format using:
       - Similar outputs from call chain (priority)
       - Simulated values (fallback)
      Set {"status": "success", "type": "mock"}
   - CASE 3: Error
      If not correct
      Set {"status": "success", "type": "error"}
2. Response Requirements:
    Strictly use JSON format only
    Never explain parameter sources or chain structure
    Never ask follow-up questions
    Maintain consistent parameter naming conventions
3. Output Format (JSON):
  "status": "<success>", // Always ’success’ per operation completion
  "data": <output_parameters>,
  "type": "<success/mock/error>"
Implementation Notes:
1. Priority Order:
   History Match > Structural Clone > Contextual Moc
API call chain is as follows:
E.3.4 Simulated User Response Agent Prompts
Simulated user response agent prompts: Utilize a parameter extractor as the user response to agent, serving as a simulated responder for follow-up questions by the agent. Strictly adhere to the parameter records of the API call chain to return only the queried and existent original parameter values. Automatically filter out uninvoked or null parameters to ensure that the responses include only the actual request information from the existing chain of calls.
As an API chain parameter extractor, directly return exact parameter values
from the given API workflows without any modification.
## Mandatory Protocols
1. Parameter Extraction Priority
    Always return raw parameter values from the latest API call
    Return empty string for blank parameters (e.g. param-cuisines_1 -> "")
2. Response Requirements
    Merge multiple parameters in single response
     Example: "patient_id:[value] cuisine:[value]"
    Strictly avoid explanations or disclaimers
    Never reveal API structure or workflow logic
## Critical Examples
User: What’s the patient ID and dietary preferences?
API Context: [param-patient_id_10:’P123’ ...]
Response: patient_id:P123’’
User: Current trial phase and calories limit?
API Context: [param-trial_phase_1:’Phase 2’ param-calories_max_1:’2000’...]
Response: phase:Phase 2 calories_max:2000
User: How to activate international roaming?
API Context: Relevant records
Response: I don’t Know international roaming activation information.
## Execution Context
Current API call chain:
E.4 Evaluation
E.4.1 Evaluation Prompts
Evaluation prompts in GPT-4.1 are designed to assess the correctness of the answer generation process, logical consistency, and accuracy of responses by analyzing the anticipated pathways and the decision-making pathways of the agent.
As an expert in response quality evaluation, you need to perform the
following steps:
I. Core Information Comparison Requirements
1. Reference Path Analysis
- Understand the simulated nature of reference API call paths.
- Be aware of potential discrepancies: API names/parameter formats may
  differ from actual implementations.
2. Actual Path Verification
- Compare each actual call path with the reference path.
- Focus on logical coherence rather than exact matching.
II. Error Detection Standards
1. Call Process Errors
 Parameter Anomalies:
  * Includes fictitious or illegal parameters.
 Execution Errors:
  * Returns error codes (e.g., 5xx) or invalid responses.
2. Information Integrity Errors
 Deviation in Answers:
  * Fails to address the core user query accurately.
 Missing Key Information:
  * Lacks necessary data items or explanation steps.
III. Correctness Determination Rules
1. Process Compliance
- Call sequence should be logically consistent.
2. Answer Completeness
- Covers all core elements of the user’s question.
- Output provides a sufficient amount of information.
IV. Quality Rating System
[1] High-Quality Standard:
* Complete logical coherence in call paths.
* Output results are accurate and effective.
* No technical errors.
[0] Deficiency Standard (if any condition is met):
* Critical API call failures.
* Returned results do not support the answer.
* Presence of unaddressed critical errors.
V. Output Specifications
1. Detection Report Format:
    1. Parameter Validation -> Compliant/Non-compliant
    2. Path Verification -> Compliant/Non-compliant
    3. Result Completeness -> Compliant/Non-compliant
2. Final Conclusion Format:
{’Quality Result’: 1} or {’Quality Result’: 0}
VI. Input Data Interface
User Question: {question}
[AGENT Answer Start]
{reference}
[AGENT Answer End]
[Reference Call Path]
{reference_chain}
[Reference Call Path End]
[Actual Call]
{agent_actual_chain}
[Actual Call End]
Appendix F Experiments on API-Bank
Model Method Easy Medium Hard All TCR TSR Steps TCR TSR Steps TCR TSR Steps TCR TSR Steps Qwen2.5-14B Decider 56.1 38.6 4.37 68.2 31.2 4.94 76.3 34.6 5.20 70.5 33.8 4.98 D+N(Alpha) 64.9 49.1 4.83 72.7 36.4 5.16 68.7 36.0 5.03 69.8 37.8 5.05 D+N(Heur) 73.7 56.1 4.88 67.0 40.9 5.25 66.8 33.6 5.46 67.8 39.4 5.27 Qwen2.5-32B Decider 68.4 50.9 4.17 84.1 43.2 5.59 72.5 36.5 6.01 76.6 41.0 5.54 D+N(Alpha) 89.5 68.4 4.47 80.1 35.8 5.70 85.3 45.0 6.03 83.8 44.4 5.62 D+N(Heur) 91.2 73.7 4.55 87.5 46.0 5.73 84.8 45.5 5.63 86.7 49.3 5.46 Deepseek-R1-32B Decider 91.2 71.9 5.33 73.3 36.4 6.44 72.5 30.8 6.42 75.2 38.3 6.16 D+N(Alpha) 89.5 63.2 5.63 90.9 48.3 6.59 81.5 34.1 6.43 86.3 43.5 6.35 D+N(Heur) 99.7 77.3 5.03 89.6 46.8 6.21 79.2 35.2 6.80 86.0 45.2 6.17 Deepseek-V3 Decider 89.5 73.7 4.80 90.9 52.3 5.47 92.4 47.9 5.71 91.4 52.9 5.45 D+N(Alpha) 99.9 82.5 4.67 98.9 58.5 6.05 99.1 51.2 5.54 99.1 58.1 5.58 D+N(Heur) 99.9 89.5 5.10 99.4 61.4 5.74 99.1 52.6 5.72 99.3 60.8 5.61 GPT-4o Decider 99.9 80.7 4.28 99.4 56.4 4.96 98.6 46.5 4.93 99.1 54.8 4.82 D+N(Alpha) 99.9 78.5 4.04 98.9 56.4 4.98 98.6 52.2 4.57 98.9 57.2 4.64 D+N(Heur) 99.9 74.3 3.75 98.3 59.1 5.02 98.1 57.9 4.90 98.4 60.5 4.77 Qwen2.5-14B(SFT) Decider 80.7 52.6 4.57 80.7 41.6 5.12 79.6 45.5 5.04 80.2 44.6 5.00 D+N(Alpha) 94.7 64.3 4.57 89.8 44.1 5.16 88.2 42.2 5.01 89.6 45.5 4.99 D+N(Heur) 93.0 72.1 4.74 89.8 48.6 5.22 87.7 45.0 5.02 89.2 49.8 5.04
Table 5: Impact of Navigator and Search Algorithms on API-Bank. Metrics are reported as in Table 2.
Appendix G Link Prediction Evaluation
Dataset APIs Nodes Edges ACC F1 AUC
ToolBench 5501 7866 24215 76.4 77.6 0.75
API-Bank 2650 6025 10255 78.4 76.1 0.71
Table 6: Tool Graph Statistics and Link Prediction Evaluation. Nodes and Edges denote the number of nodes and edges in the graph, respectively. ACC and F1 are reported as percentages (%), while AUC is reported as a value between 0 and 1.